OpenStudy (anonymous):

The sum of two consecutive integers is between 5 and 9. What are the two integers? Tell the two numbers only.

Parth (parthkohli):

No, 6 + 7 = 13

OpenStudy (anonymous):

3 and 4

OpenStudy (anonymous):

that is what i said didn't want to give away answer

Parth (parthkohli):

2,3; 3,4; 4,5

OpenStudy (anonymous):

3,4

OpenStudy (anonymous):

Medal

Parth (parthkohli):

2+3 = 5, 3+4 = 7 and 4+5 = 9... Now it's between, so excluding... Answer is 3 and 4

Parth (parthkohli):

maddog, you have to explain too...

OpenStudy (anonymous):

ok i will medal